OpenStudy (anonymous):

Simplify: [sqrt(2)^sqrt(3)][sqrt(2)^sqrt(3)]

OpenStudy (anonymous):

\[\sqrt{2}^{\sqrt{3}}\sqrt{2}^{\sqrt{3}}\]

OpenStudy (anonymous):

\[\sqrt{2}^{2\sqrt{3}}=\left( \sqrt{2}^2 \right)^\sqrt{3}=2^\sqrt{3}\]from here you can only approximate the answer using a calculator

OpenStudy (anonymous):

k thanks!